OpenStudy (anonymous):

Given right triangle ABC right angle at C, if A=5 and B=12 use pythagorean theorem to solve for C

OpenStudy (anonymous):

Pythagorean theorem is \(C=\sqrt{B^2+A^2}\)

OpenStudy (texaschic101):

A^2 + B^2 = C^2 5^2 + 12^2 = C^2 25 + 144 = C^2 169 = C^2 sq rt 169 = C 13 = C any questions ?

OpenStudy (anonymous):

Thank you! and no that pretty much explained it.